In one town, the number of burglaries in a week has a Poisson distribution with mean μ = 3.5. Let variable x denote the number of burglaries in this town in a randomly selected month. Find the smallest usual value for x. Round your answer to three decimal places.

(HINT: Assume a month to be exactly 4 weeks)
Answer #1

Smallest usual value = mean - 2*(standard deviation)

We need it for month hence:

Smallest usual value = 14 - 2*3.7417 = 6.5166

Hence the smallest usual value for x is 6.517 (rounded to 3 decimal places)

Please comment if any doubt. Thank you.